# Characterizing Kadison–Schwarz maps on $M_{3}$

### v1

@@ Abstract
Kadison–Schwarz (KS) maps form a natural class of positive linear maps lying between positivity and complete positivity. Despite their relevance in quantum dynamics and operator algebras, a detailed characterization of KS maps beyond low dimensions remains largely open. In this work we analyze unital linear maps on $M_{3}$ using the Bloch–Gell–Mann representation. Exploiting unitary equivalence and structural properties of the $\mathfrak{su}(3)$ algebra, we derive analytic conditions ensuring the Kadison–Schwarz property. Our approach clarifies the relation between KS maps and completely positive maps on $M_{3}$.
@@ II Kadison–Schwarz maps
We next recall several basic properties of unital positive maps that will be useful in the analysis of the Kadison–Schwarz condition, starting with a classical result due to Kadison.
@@
The following lemma clarifies to what extent the KS inequality is automatically satisfied by unital positive maps.
Lemma 2.
Let $\Phi:M_{d}\to M_{d}$ be a unital positive map. Then the Kadison–Schwarz inequality (1) holds for all normal operators $X\in M_{d}$, i.e., for operators satisfying $X^{\dagger}X=XX^{\dagger}$.
Proof.
Let $X$ be a normal operator. Then $X^{\dagger}X=XX^{\dagger}\geq 0$ and $\|X\|^{2}=\|X^{\dagger}X\|$. By positivity, $\Phi(X^{\dagger}X)\geq 0$. Using Lemma 1, we obtain
$\|\Phi(X)\|^{2}\;\leq\;\|X\|^{2}\;=\;\|X^{\dagger}X\|.$
For any operator $B$, one has $B^{\dagger}B\leq\|B\|^{2}\mathbb{I}$, and hence
$\Phi(X)^{\dagger}\Phi(X)\;\leq\;\|\Phi(X)\|^{2}\mathbb{I}\;\leq\;\|X^{\dagger}X\|\,\mathbb{I}.$
Since both $\Phi(X^{\dagger}X)$ and $\Phi(X^{\dagger})\Phi(X)$ are positive operators bounded above by the same scalar multiple of the identity, the inequality (1) follows. ∎
Lemma 2 shows that the KS property does not impose additional constraints on unital positive maps when restricted to normal operators. Consequently, possible violations of the KS inequality must originate from non-normal operators, cf. [Chruscinski2010, Mukhamedov2011].

### v2

@@ Abstract
Kadison–Schwarz (KS) maps form a natural class of positive linear maps lying between positivity and complete positivity. Despite their relevance in quantum dynamics and operator algebras, a detailed analytic characterization of KS maps is still largely lacking. In this work we analyze unital linear maps on $M_{3}$ using the Bloch–Gell–Mann representation. Exploiting unitary equivalence and structural properties of the $\mathfrak{su}(3)$ algebra, we derive explicit analytic conditions ensuring the Kadison–Schwarz property. Our approach clarifies the relation between KS maps and completely positive maps on $M_{3}$.
@@ II Kadison–Schwarz maps
We next recall a basic property of unital positive maps that will be useful in the analysis of the Kadison–Schwarz condition.
@@
It is worth noting that for unital positive maps the Kadison–Schwarz inequality is automatically satisfied when restricted to normal operators, whereas possible violations of the KS condition necessarily involve non-normal elements [Chruscinski2010, Mukhamedov2011].
@@ V Conclusions
Acknowledgements
The author would like to thank Aristides Katavolos for helpful comments that improved the clarity of the manuscript.